7. Consider the generalized Pell equation x2- 6y2 =3. *) Verify that 3+ v6 E Z[V6] gives a solution of (*). Produce three othplease help!
Answer #1

Clearly, 3+√6 is a solution of x² - 6y² = 3 as, 3² - 6.1² = 3

Note the following theorem :

Fix u = a + bVd where a2 - db2 Theorem = 1 with a and b in Z+. For each {0}, every solution of x2 - dy2 Jaæ VInu and |y| < V[

We will use the results of this theorem in the following main proof.

We will describe all the solutions of x2 - 6y2 solution is (3, 1) and well show every solution is a Pell multiple of (3,1) p